1. Factor completely:
x3y – 3x2y – 10xy

2. Factor completely:
x2 – 4xy + 12ax – 48ay

3. Factor completely:
x2 – 5x – 6

4. Factor completely:
36x2 – 4y2

5. Factor completely:
9x2 + 40x + 16

6. Factor completely:
3x2 – 30x + 48

can you please help me on these? thank you

Answer by stanbon(57342)

1. Factor completely:
x^3y – 3x^2y – 10xy
=xy(x^2-3x-10)
=xy(x-5)(x+2)
----------------------
2. Factor completely:
x^2 – 4xy + 12ax – 48ay
=x(x-4y)+12a(x-4y)
=(x-4y)(x+12a)
-----------------
3. Factor completely:
x^2 – 5x – 6
=(x-6)(x+1)
----------------------------------
4. Factor completely:
36x^2 – 4y^2
=4(9x^2-y^2)
=4(3x-y)(3x+y)
----------------------------
5. Factor completely:
9x2 + 40x + 16
=(9x+4)(x+4)
---------------
6. Factor completely:
3x^2 – 30x + 48
=3(x^2-10x+16)
=3(x-8)(x-2)
-----------------
Cheers,
Stan H.